# The long way of a viscous vortex dipole

### v1

@@ 4.3. Bound on the vertical speed
Before setting up the nonlinear energy estimates, we apply Lemmas 4.54.6 to control the size of the correction $\zeta$ to the vertical speed.
@@
To bound the term involving $\Omega_{\mathrm{app}}$, we split the integration domain into the region $\mathrm{I}_{\varepsilon}$ and its complement $\mathrm{I}_{\varepsilon}^{c}$, and we apply Hölder’s inequality with $11/p+1/q=$ 1 and $q>2$. We thus find
$|\langle\partial_{1}\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}\rangle_{L^{2}}|\,\lesssim\,\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-3}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\|(1+|\cdot|)^{3}\Omega_{\mathrm{app}}\|_{L^{p}}+\|\partial_{1}\varphi\|_{L^{q}}\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}\Omega_{\mathrm{app}}\|_{L^{p}}\,.$
As $\Omega_{\mathrm{app}}\in\mathcal{Z}$ and $\mathrm{I}_{\varepsilon}^{c}\subset\{\xi\in\mathbb{R}^{2}\,;\,|\xi|>\varepsilon^{-\sigma_{1}}\}$ by Proposition 4.4, it is straightforward to verify that $\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}\Omega_{\mathrm{app}}\|_{L^{p}}\leq\exp(-c_{*}\varepsilon^{-2\sigma_{1}})$ for some $c_{*}>0$. Therefore using Lemma 4.6 to bound the first term in the right-hand side and Lemma 4.5 for the second one, we arrive that
@@
$|\langle\partial_{1}\mathcal{T}_{\varepsilon}\varphi,w\rangle_{L^{2}}|\,\lesssim\,\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-3}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\|(1+|\cdot|)^{3}w\|_{L^{p}}+\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}W_{\varepsilon}^{-1/2}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{2}}\|w\|_{\mathcal{X}_{\varepsilon}}\,.$
Since $21<p<$ 2, Hölder’s inequality readily implies that $\|(1+|\cdot|)^{3}w\|_{L^{p}}\lesssim\|w\|_{\mathcal{X}_{\varepsilon}}$. Thus, invoking Lemmas 4.5 and 4.6, and using again Hölder’s inequality with $21/q+1/r=1/$ 2, we obtain
@@ 4.4. The energy functional
Following [13, Section 4.3], the idea is to compare $E_{\varepsilon}$ with $E_{0}$. First of all, using Lemma 4.6 and arguing as in the proof of (4.39), it is not difficult to verify that
@@
$\begin{split}E_{\varepsilon}^{1}[w]\,&=\,E_{\varepsilon}^{1}[w_{\mathrm{in}}]+E_{\varepsilon}^{1}[w_{\mathrm{out}}]+\frac{1}{2}\langle\varphi_{\mathrm{in}},w_{\mathrm{out}}\rangle_{L^{2}}+\frac{1}{2}\langle\varphi_{\mathrm{out}},w_{\mathrm{in}}\rangle_{L^{2}}\\ \,&=\,E_{\varepsilon}^{1}[w_{\mathrm{in}}]+E_{\varepsilon}^{1}[w_{\mathrm{out}}]+\langle\varphi_{\mathrm{in}},w_{\mathrm{out}}\rangle_{L^{2}}\,,\end{split}$ (4.45)
where in the second line we integrate by parts using $w_{\mathrm{in}}=\Delta\varphi_{\mathrm{in}}$ and $w_{\mathrm{out}}=\Delta\varphi_{\mathrm{out}}$. To estimate the last term in (4.45), we apply Hölder’s inequality with $11/q+1/p=$ 1 and $q>2$, $21<p<$ 2. Using (4.31) to bound the stream function $\varphi_{\mathrm{in}}$ and recalling that $w_{\mathrm{out}}$ is supported outside the region $\mathrm{I}_{\varepsilon}$, we obtain
@@
Finally, combining (4.47), (4.49) and (4.51), we obtain
@@ 4.5. The energy identity
as long as $t<T_{\mathsf{adv}}\delta^{-\sigma}$. As a consequence, the energy (4.40) satisfies the estimate
@@
In view of Propositions 3.2 and 4.8, estimate (4.61) implies that the solution $w$ of (4.4)–(4.6) does not become much larger than the source term $\delta^{-2}\mathcal{R}_{M}$, which is of size $\delta^{-2}\varepsilon^{M+1}+\varepsilon^{2}$.
@@
The constant $s_{*}$ in (4.58)–(4.60) can be taken as a small multiple of $s_{0}$.
@@ 4.6. The diffusion functional
$-\frac{1}{2}\int_{\mathrm{II}_{\varepsilon}}\bigl{(}t\partial_{t}(\exp(\varepsilon(t)^{-2\sigma_{1}}/4))\bigr{)}|w|^{2}\,{\rm d}\xi\,=\,\frac{\sigma_{1}}{8}\int_{\mathrm{II}_{\varepsilon}}\rho_{\varepsilon}^{2}|w|^{2}W_{\varepsilon}\,{\rm d}\xi\,=\,\frac{\sigma_{1}}{8}\left\lVert\mathbbm{1}_{\mathrm{II}_{\varepsilon}}\rho_{\varepsilon}w\right\rVert^{2}_{\mathcal{X}_{\varepsilon}}\,.$ (4.68)
@@
We also observe that $\langle\nabla w,\nabla\varphi\rangle_{L^{2}}=-\|w\|_{L^{2}}^{2}$ and $\langle w,\xi\cdot\nabla\varphi\rangle_{L^{2}}=0$. The first equality is a simple integration by parts, and the second one is conveniently obtained by using polar coordinates $r,\theta$ defined by $\xi_{1}=r\cos\theta$, $\xi_{2}=r\sin\theta$. Indeed, since both operators $r^{-1}\partial_{r}$ and $\partial_{\theta}$ are (at least formally) skew-symmetric in the space $L^{2}_{r}:=L^{2}((0,+\infty)\times(0,2\pi)\,,r{\rm d}r{\rm d}\theta)$, we have
[…]
@@
Step 2: Decomposition of the diffusive quadratic form. The expression (4.74) is the analogue of the quadratic form $Q_{\varepsilon}[w]$ appearing in [14, Section 4.7]. Taking formally the limit $\varepsilon\to 0$, we obtain
@@
which is exactly the quadratic form $Q_{0}[w]$ studied [13, 14]. In particular, it is established in [14, Proposition 4.14] that, for any $w\in\mathcal{X}_{0}$ with $\rho_{0}w\in\mathcal{X}_{0}$ and $\nabla w\in\mathcal{X}_{0}^{2}$, the following holds
@@
$\begin{split}\mathcal{D}_{\mathcal{L},\varepsilon}[w]\,&=\,\mathcal{D}_{\mathcal{L},\varepsilon}[w_{\mathrm{in}}]+\mathcal{D}_{\mathcal{L},\varepsilon}[w_{\mathrm{out}}]-\left\lVert\sqrt{|\nabla\chi_{1}|^{2}+|\nabla\chi_{2}|^{2}}\,w\right\rVert_{\mathcal{X}_{\varepsilon}}^{2}\,,\\ \,&\geq\,\mathcal{D}_{\mathcal{L},\varepsilon}[w_{\mathrm{in}}]+\mathcal{D}_{\mathcal{L},\varepsilon}[w_{\mathrm{out}}]-C_{4}\varepsilon^{2\sigma_{1}}\|w\|_{\mathcal{X}_{\varepsilon}}^{2}\,,\end{split}$ (4.79)
@@
$\begin{split}\mathcal{I}_{\varepsilon}\bigl{[}\mathbbm{1}_{\mathrm{I}_{\varepsilon}}w_{\mathrm{out}}\bigr{]}\,&\geq\,\frac{1}{32}\,\bigl{\|}\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\rho_{0}w_{\mathrm{out}}\bigr{\|}_{\mathcal{X}_{0}}^{2}-C\bigl{(}\varepsilon^{2\sigma_{1}}+\varepsilon^{\gamma_{2}+\sigma_{1}}\bigr{)}\left\lVert\rho_{\varepsilon}w\right\rVert_{\mathcal{X}_{\varepsilon}}^{2}\\ \,&\geq\,\frac{1}{32}\,\bigl{\|}\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\rho_{\varepsilon}w_{\mathrm{out}}\bigr{\|}_{\mathcal{X}_{\varepsilon}}^{2}-C_{6}\bigl{(}\varepsilon^{2\sigma_{1}}+\varepsilon^{\gamma_{2}}\bigr{)}\left\lVert\rho_{\varepsilon}w\right\rVert_{\mathcal{X}_{\varepsilon}}^{2}\,,\end{split}$ (4.88)
@@ 4.7. The advection terms
In this section we estimate the advection terms defined in (4.54), which are potentially dangerous because of the large prefactor $\delta^{-1}$. In the inner region $\mathrm{I}_{\varepsilon}$, we exploit crucial cancellations related to the structure of the energy functional (4.40), which were explained in an informal way in Section 4.1. In the other regions the dominant term in the energy is the weighted enstrophy $\frac{1}{2}\|w\|_{\mathcal{X}_{\varepsilon}}^{2}$, and the influence of the large advection terms can be controlled by an appropriate choice of the weight function $W_{\varepsilon}$.
@@
$\mathrm{A}_{1}\,$ =$=\,\left\langle\left\{\Phi_{\mathrm{app}}^{E},w\right\},W_{\varepsilon}w\right\rangle_{L^{2}}\,,$ (4.90)
$\mathrm{A}_{2}\,$ =$=\,\left\langle\left\{\varphi-\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}^{E}\right\},W_{\varepsilon}w\right\rangle_{L^{2}}+\left\langle\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\left\{\Phi_{\mathrm{app}}^{E},w\right\},\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}$ (4.91)
@@
$\mathrm{A}_{NS}\,$ =$=\,\left\langle\left\{\Phi_{\mathrm{app}}^{NS},w\right\}+\left\{\varphi-\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}^{NS}\right\},W_{\varepsilon}w+\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}\,.$ (4.93)
@@
$|\mathrm{A}_{NS}|\,$ $\lesssim\,\varepsilon\bigl{(}\|w\|_{\mathcal{X}_{\varepsilon}}^{2}+\|\nabla w\|_{\mathcal{X}_{\varepsilon}}^{2}\bigr{)}\,,$ (4.98)
@@
$\mathrm{A}_{1}\,$ =$=\,\left\langle\left\{\Phi_{\mathrm{app}}^{E},w\right\},W_{\varepsilon}w\right\rangle_{L^{2}}\,=\,-\frac{1}{2}\left\langle\left\{\Phi_{\mathrm{app}}^{E},W_{\varepsilon}\right\}w,w\right\rangle_{L^{2}}$
@@
To control the first term in (4.99), we use the bound (4.21) on $\nabla\Theta$ and the estimate (4.26) on $\nabla W_{\varepsilon}$ in region $\mathrm{I}_{\varepsilon}$, which give |$|\{\Theta,W_{\varepsilon}\}|\lesssim\varepsilon^{M+1}(1+|\xi|)^{N}W_{\varepsilon}$ for some $N\in\mathbb{N}$. Since $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$ in that region, we infer that
@@
[…]
@@
$|\left\langle\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\left\{\Theta,w\right\},\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}|\,\lesssim\,\varepsilon^{M+1}\|w_{\varepsilon}\|_{\mathcal{X}_{\varepsilon}}\|\nabla w\|_{\mathcal{X}_{\varepsilon}}\,.$ (4.103)
@@
$\begin{split}&|\left\langle\mathbbm{1}_{\mathrm{I}\mathrm{I}_{\varepsilon}\cup\mathrm{III}_{\varepsilon}}\{w,\Omega_{\mathrm{app}}^{E}\}W_{\varepsilon}+\mathbbm{1}_{\mathrm{I}\mathrm{I}\mathrm{I}_{\varepsilon}}\left\{W_{\varepsilon},\Omega_{\mathrm{app}}^{E}\right\}w,\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}|\\ &\qquad\qquad\qquad\,\lesssim\,\exp(-c_{*}\varepsilon^{-2\sigma_{1}})\|w\|_{\mathcal{X}_{\varepsilon}}(\|w\|_{\mathcal{X}_{\varepsilon}}+\|\nabla w\|_{\mathcal{X}_{\varepsilon}})\,,\end{split}$ (4.104)
@@
[…]
|$\bigl{|}\left\langle\left\{\Phi_{\mathrm{app}}^{NS},w\right\},W_{\varepsilon}w+\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}\bigr{|}\,\lesssim\,\varepsilon\,\|\nabla w\|_{\mathcal{X}_{\varepsilon}}\|w\|_{\mathcal{X}_{\varepsilon}}\,,$ (4.106)
@@
Finally, if we assume that $\sigma_{1}>0$ is small enough so that $\sigma_{1}N\leq(M+1)/2$, and $\sigma_{2}>1$ large enough so that $21+\sigma_{2}\geq(M+1)/$ 2, it follows from (4.63) that
@@
provided $40<s_{*}\leq s_{0}(M+1)/$ 4. Under this assumption the bound (4.58) is a straightforward consequence of (4.95)–(4.98) and Proposition 4.12. ∎
4.8. The forcing term

### v2

@@ 4.3. Bound on the vertical speed
Before setting up the nonlinear energy estimates, we apply Lemma 4.5 and Corollary 4.7 to control the size of the correction $\zeta$ to the vertical speed.
@@
To bound the term involving $\Omega_{\mathrm{app}}$, we split the integration domain into the region $\mathrm{I}_{\varepsilon}$ and its complement $\mathrm{I}_{\varepsilon}^{c}$, and we apply Hölder’s inequality with $1/p+1/q=1$ and $q>2$. We thus find
$|\langle\partial_{1}\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}\rangle_{L^{2}}|\,\lesssim\,\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-1}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\|(1+|\cdot|)\Omega_{\mathrm{app}}\|_{L^{p}}+\|\partial_{1}\varphi\|_{L^{q}}\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}\Omega_{\mathrm{app}}\|_{L^{p}}\,.$
As $\Omega_{\mathrm{app}}\in\mathcal{Z}$ and $\mathrm{I}_{\varepsilon}^{c}\subset\{\xi\in\mathbb{R}^{2}\,;\,|\xi|>\varepsilon^{-\sigma_{1}}\}$ by Proposition 4.4, it is straightforward to verify that $\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}\Omega_{\mathrm{app}}\|_{L^{p}}\leq\exp(-c_{*}\varepsilon^{-2\sigma_{1}})$ for some $c_{*}>0$. Therefore using Corollary 4.7 to bound the first term in the right-hand side and Lemma 4.5 for the second one, we arrive that
@@
$|\langle\partial_{1}\mathcal{T}_{\varepsilon}\varphi,w\rangle_{L^{2}}|\,\lesssim\,\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-1}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{q}}\|(1+|\cdot|)w\|_{L^{p}}+\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}^{c}}W_{\varepsilon}^{-1/2}\partial_{1}\mathcal{T}_{\varepsilon}\varphi\|_{L^{2}}\|w\|_{\mathcal{X}_{\varepsilon}}\,.$
Since $1<p<2$, Hölder’s inequality readily implies that $\|(1+|\cdot|)w\|_{L^{p}}\lesssim\|w\|_{\mathcal{X}_{\varepsilon}}$. Thus, using again (4.31), (4.35) and applying Hölder’s inequality with $1/q+1/r=1/2$, we obtain
@@
Following [14, Section 4.3], the idea is to compare $E_{\varepsilon}$ with $E_{0}$. First of all, using Corollary 4.7 and arguing as in the proof of (4.40), it is not difficult to verify that
@@
$w\,=\,\mathbbm{1}_{\mathrm{I}_{\varepsilon}}w+\bigl(1-\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\bigr)w\,=:\,w_{\mathrm{in}}+w_{\mathrm{out}}\,.$
@@
where in the second line we integrate by parts using $w_{\mathrm{in}}=\Delta\varphi_{\mathrm{in}}$ and $w_{\mathrm{out}}=\Delta\varphi_{\mathrm{out}}$. To estimate the last term in (4.46), we apply Hölder’s inequality with $1/q+1/p=1$ and $q>2$, $1<p<2$. Using (4.31) to bound the stream function $\varphi_{\mathrm{in}}$ and recalling that $w_{\mathrm{out}}$ is supported outside the region $\mathrm{I}_{\varepsilon}$, we obtain
@@
Finally, combining (4.48), (4.49), (4.50) and (4.52), we obtain
@@ 4.5. The energy identity
as long as $t\leq T_{\mathsf{adv}}\delta^{-\sigma}$. As a consequence, the energy (4.41) satisfies the estimate
@@
In view of Propositions 3.2 and 4.9, estimate (4.62) implies that the solution $w$ of (4.4)–(4.6) does not become much larger than the source term $\delta^{-2}\mathcal{R}_{M}$, which is of size $\delta^{-2}\varepsilon^{M+1}+\varepsilon^{2}$.
@@
The constant $s_{*}$ in (4.59), (4.61) can be taken as a small multiple of $s_{0}$.
@@ 4.6. The diffusion functional
$\frac{1}{2}\int_{\mathrm{II}_{\varepsilon}}\bigl(t\partial_{t}(\exp(\varepsilon(t)^{-2\sigma_{1}}/4))\bigr)|w|^{2}\,{\rm d}\xi\,=\,\frac{\sigma_{1}}{8}\int_{\mathrm{II}_{\varepsilon}}\rho_{\varepsilon}^{2}|w|^{2}W_{\varepsilon}\,{\rm d}\xi\,=\,\frac{\sigma_{1}}{8}\left\lVert\mathbbm{1}_{\mathrm{II}_{\varepsilon}}\rho_{\varepsilon}w\right\rVert^{2}_{\mathcal{X}_{\varepsilon}}\,.$ (4.69)
@@
Similarly, since $w=\Delta\varphi$, we observe that $\langle\nabla w,\nabla\varphi\rangle_{L^{2}}=-\|w\|_{L^{2}}^{2}$ and
$\langle w,\xi\cdot\nabla\varphi\rangle_{L^{2}}\,=\,\langle\Delta\varphi,\xi\cdot\nabla\varphi\rangle_{L^{2}}\,=\,-\int_{\mathbb{R}^{2}}\Bigl(|\nabla\varphi|^{2}+\frac{1}{2}\,\xi\cdot\nabla\,|\nabla\varphi|^{2}\Bigr){\rm d}\xi\,=\,0\,.$
@@
Step 2: Decomposition of the diffusive quadratic form. The expression (4.74) is the analogue of the quadratic form $Q_{\varepsilon}[w]$ appearing in [15, Section 4.7]. Taking formally the limit $\varepsilon\to 0$, we obtain
@@
which is exactly the quadratic form $Q_{0}[w]$ studied [14, 15]. In particular, it is established in [15, Proposition 4.14] that, for any $w\in\mathcal{X}_{0}$ with $\rho_{0}w\in\mathcal{X}_{0}$ and $\nabla w\in\mathcal{X}_{0}^{2}$, the following holds
@@
[…]
@@
[…]
@@ 4.7. The advection terms
In this section we estimate the advection terms defined in (4.55), which are potentially dangerous because of the large prefactor $\delta^{-1}$. In the inner region $\mathrm{I}_{\varepsilon}$, we exploit crucial cancellations related to the structure of the energy functional (4.41), which were explained in an informal way in Section 4.1. In the other regions the dominant term in the energy is the weighted enstrophy $\frac{1}{2}\|w\|_{\mathcal{X}_{\varepsilon}}^{2}$, and the influence of the large advection terms can be controlled by an appropriate choice of the parameters (4.19).
@@
$\mathrm{A}_{1}\,$ $=\,\left\langle\left\{\Phi_{\mathrm{app}}^{E},w\right\},W_{\varepsilon}w\right\rangle_{L^{2}}\,$, (4.90)
$\mathrm{A}_{2}\,$ $=\,\left\langle\left\{\varphi-\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}^{E}\right\},W_{\varepsilon}w\right\rangle_{L^{2}}+\left\langle\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\left\{\Phi_{\mathrm{app}}^{E},w\right\},\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}$ (4.91)
@@
$\mathrm{A}_{NS}\,$ $=\,\left\langle\left\{\Phi_{\mathrm{app}}^{NS},w\right\}+\left\{\varphi-\mathcal{T}_{\varepsilon}\varphi,\Omega_{\mathrm{app}}^{NS}\right\},W_{\varepsilon}w+\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}\,.$ (4.93)
@@
$|\mathrm{A}_{NS}|\,$ $\lesssim\,\varepsilon^{2}\bigl(\|w\|_{\mathcal{X}_{\varepsilon}}^{2}+\|\nabla w\|_{\mathcal{X}_{\varepsilon}}^{2}\bigr)\,$, (4.98)
@@
$\mathrm{A}_{1}\,$ $=\,\left\langle\left\{\Phi_{\mathrm{app}}^{E},w\right\},W_{\varepsilon}w\right\rangle_{L^{2}}\,=\,-\frac{1}{2}\left\langle\left\{\Phi_{\mathrm{app}}^{E},W_{\varepsilon}\right\}w,w\right\rangle_{L^{2}}$
@@
To control the first term in (4.99), we use the bound (4.21) on $\nabla\Theta$ and the estimate (4.26) on $\nabla W_{\varepsilon}$ in region $\mathrm{I}_{\varepsilon}$, which give $|\{\Theta,W_{\varepsilon}\}|\lesssim\varepsilon^{M+1}(1+|\xi|)^{N}W_{\varepsilon}$ for some $N\in\mathbb{N}$. Since $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$ in that region, we infer that
@@
$\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\bigl\{\Omega_{\mathrm{app}}^{E},w\bigr\}W_{\varepsilon}\,=\,\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\bigl\{F(\Omega_{\mathrm{app}}^{E}),w\bigr\}\,=\,\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\bigl\{\Theta,w\bigr\}-\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\bigl\{\Phi_{\mathrm{app}}^{E},w\bigr\}\,$,
@@
$|\left\langle\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\left\{\Theta,w\right\},\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}|\,\lesssim\,\varepsilon^{M+1}\|w\|_{\mathcal{X}_{\varepsilon}}\|\nabla w\|_{\mathcal{X}_{\varepsilon}}\,.$ (4.103)
@@
[…]
@@
On the other hand, using the definition (4.9), the bound (4.32), and the fact that $\Omega_{\mathrm{app}}^{NS}=\mathcal{O}_{\mathcal{Z}}(\varepsilon^{2})$ and $\zeta_{\mathrm{app}}^{NS}=\mathcal{O}(\varepsilon)$, one verifies that $\|\nabla\Phi_{\mathrm{app}}^{NS}\|_{L^{\infty}}\lesssim\varepsilon^{2}$. This gives
$\bigl|\left\langle\left\{\Phi_{\mathrm{app}}^{NS},w\right\},W_{\varepsilon}w+\varphi-\mathcal{T}_{\varepsilon}\varphi\right\rangle_{L^{2}}\bigr|\,\lesssim\,\varepsilon^{2}\,\|\nabla w\|_{\mathcal{X}_{\varepsilon}}\|w\|_{\mathcal{X}_{\varepsilon}}\,$, (4.106)
@@
Finally, if we assume that $\sigma_{1}>0$ is small enough so that $\sigma_{1}N\leq(M+1)/2$, and $\sigma_{2}>1$ large enough so that $1+\sigma_{2}\geq(M+1)/2$, it follows from (4.64) that
@@
provided $0<s_{*}\leq s_{0}(M+1)/4$. Under this assumption the bound (4.59) is a straightforward consequence of (4.95)–(4.98) and Proposition 4.13. ∎
4.8. The forcing terms